OpenStudy (nickthurmm1):

Isosceles trapezoid ABCD has legs If AB = 4y – 3, BC = 3y – 4, and CD = 5y – 10, find the value of y.

OpenStudy (squidnnneyyy):

is there a diagram?

OpenStudy (dragonman1):

or a picture that you can attach

OpenStudy (nickthurmm1):

no

OpenStudy (squidnnneyyy):

All you have to do is imagine that you have a trapezoid in front of you that is labeled ABCD (A is the top left, B is the top right, C is the bottom right, and D is the bottom left.)

OpenStudy (squidnnneyyy):

Then imagine the diagonals being connected by two lines that form an x.

OpenStudy (squidnnneyyy):

I would draw this out soyou see what I mean

OpenStudy (squidnnneyyy):

AD is equal to BC

OpenStudy (dragonman1):

so ad and bc would be the sides right

OpenStudy (squidnnneyyy):

yes

OpenStudy (squidnnneyyy):

they are equal becase this is an isosceles trapezoid so two sides are equal

OpenStudy (dragonman1):

so it would look like this

OpenStudy (squidnnneyyy):

move everything clockwise 1, so B would be A, C would be Betc...

OpenStudy (squidnnneyyy):

got it? Want me to check again, or keep explaining?

OpenStudy (dragonman1):

ye i got it

OpenStudy (squidnnneyyy):

ok so then you would use the formula AB=CD subtsitue in the equation for each. 4y-3=5y-10 Then solve. -4y Subtract 4y on each side

OpenStudy (squidnnneyyy):

-3=y-10 +10 Add 10 on each side 7=y

OpenStudy (dragonman1):

yes good job

OpenStudy (squidnnneyyy):

Thanks!